Use big ๐‘€ method solve. ๐‘€๐‘–๐‘›๐‘–๐‘š๐‘–๐‘ง๐‘’ ๐‘ = 2๐‘ฆ1 + 4๐‘ฆ2 ๐‘†๐‘ข๐‘๐‘—๐‘’๐‘๐‘ก ๐‘ก๐‘œ 2๐‘ฆ1 โ€“ 3๐‘ฆ2 โ‰ฅ 2,



โˆ’๐‘ฆ1 + ๐‘ฆ2 โ‰ฅ 3; ๐‘ฆ1



, ๐‘ฆ2 โ‰ฅ 0

1
Expert's answer
2021-11-29T17:00:17-0500

๐‘€๐‘–๐‘›๐‘–๐‘š๐‘–๐‘ง๐‘’ "\ud835\udc4d = 2\ud835\udc66_1 + 4\ud835\udc66_2"

๐‘†๐‘ข๐‘๐‘—๐‘’๐‘๐‘ก ๐‘ก๐‘œ

"2\ud835\udc66_1 \u2013 3\ud835\udc66_2 \u2265 2"

"\u2212\ud835\udc66_1 + \ud835\udc66_2 \u2265 3"

"\ud835\udc66_1, \ud835\udc66_2 \u2265 0"


After introducing surplus,artificial variables

Minย Z=2x1+4x2+0S1+0S2+MA1+MA2

subject to

2x1-3x2-S1+A1=2-x1+x2-S2+A2=3

andย x1,x2,S1,S2,A1,A2โ‰ฅ0



Positive maximumย Zj-Cjย isย M-2ย and its column index isย 1. So,ย the entering variable isย x1.

Minimum ratio isย 1ย and its row index isย 1. So,ย the leaving basis variable isย A1.

โˆดย The pivot element isย 2.

Enteringย =x1, Departingย =A1, Key Elementย =2



Since allย Zj-Cjโ‰ค0

Hence, optimal solution is arrived with value of variables as :

x1=1,x2=0

Minย Z=2

But this solution is not feasible

because the final solution violates theย 2ndย constraintย -ย ย x1ย +ย ย x2ย โ‰ฅย 3.

and the artificial variableย A2ย appears in the basis with positive valueย 4